# SonicSieve: Bringing Directional Speech Extraction to Smartphones Using Acoustic Microstructures

### v1

@@ 4.2. Subjective user evaluation of audio quality
While objective metrics like SI-SDR provide valuable quantitative assessment, they cannot fully capture the perceptual quality of speech enhancement systems in real-world scenarios. To evaluate the actual listening experience, we conduct a user study to complement our objective measurements. We randomly select 12 generated audio samples from our test set collected from Room A. The selected audio contains one target speech and 1-2 interfering sources (undesired speech or other sound). Each sample is processed by three systems: the 2-channel microphone array (baseline), 5-channel microphone array, and our proposed SonicSieve. 20 participants (14 males and 6 females) with an average age of 24 years were recruited for the listening test. The evaluation protocol consisted of three steps: (1) listening to the noisy audio mixture, (2) listening to the clean target speaker audio as reference, and (3) listening to the three processed outputs presented in random order to avoid position bias. Participants were asked to rate each processed audio sample on two criteria using 5-point scales:
@@
4.3. User Interface
@@
Finally, we consider how users can interact with their smartphone to select the desired direction. We designed and evaluated three user interfaces on Android (Fig.  12a). All interfaces display the selected direction on a semicircle. Users can make directional selections using 1) Radial Selectors by directly selecting the desired sector on the semicircle, 2) Toggle Switches by toggling switches labeled with angular ranges (e.g. $0-30^{\circ}$, $30-60^{\circ}$ etc.), or a 3) Linear Bar by selecting buttons arranged linearly across the screen and labeled with angular ranges. We also investigated how sector granularity affects usability by evaluating configurations with 6 sectors ($30^{\circ}$ each) and 9 sectors ($20^{\circ}$ each) for all three UIs.
We conducted a user study across 10 participants in a conference room, with one participant and two experimenters (Fig.  12b). For each interface, the participant completed three directional selection tasks: specifically they were tasked with selecting the first experimenter (A), the second experimenter (B), and both experimenters (A+B). During each of the three tasks, experimenters positioned themselves at different predetermined locations. The positions of the experimenters were consistent across all UI evaluations to ensure fair comparisons could be made. However, within each UI the order of tasks was randomized to minimize learning effects. Participants were instructed to select any number of sectors they felt would be needed to include the target speaker(s). We evaluated interface usability through three measurements:
Selection Accuracy. We assessed whether users correctly included the target speaker(s) (Fig.  12c). In the 6-sector condition, Toggle Switches and Linear Bar performed equally well (93.3%), while Radial Selectors exhibited the lowest accuracy (83.3%). For 9-sector condition, Linear Bar (93.3%) outperformed Radial Selectors (90.0%), with Toggle Switches showing the worst performance (86.7%). These results show that Linear Bar is the most effective interface in terms of selection accuracy, likely because it provides users with strong spatial references as well as precise numerical markers.
Interface Preference. We asked participants to rate their likelihood of using each UI on a 5-point scale (1 = extremely unlikely, 5 = extremely likely) (Fig.  12d). Participants rated Radial Selectors highest in both 6- and 9-sector settings (4.8 ± 0.6 and 4.4 ± 1.0), followed by Linear Bar (3.5 ± 1.5 and 3.1 ± 1.3), with Toggle Switches rated lowest (2.3 ± 0.8 and 2.1 ± 1.4). Users preferred Radial Selectors for its intuitive spatial mapping, while Toggle Switches’ lack of spatial correspondence led to repeated selection attempts and lower scores. Ratings were consistently higher for 6-sector interfaces, which users found easier and less error-prone to accidental selection of adjacent sectors. While most users prioritized ease of use, a few preferred Toggle Switches or Linear Bar in the 9-sector setting as it gave them greater sense precision in their selections.
Selection Confidence. We assessed users’ confidence in their selections under two conditions: when they were restricted to selecting only one sector versus when they could select multiple sectors (Fig.  12e). We performed a within-subjects experiments where we requested users to rate their selection confidence between 1 and 5 and allowed them to use their preferred UI. In the 9-sector configuration, confidence decreased from 4.3 ± 1.2 when multiple sectors were allowed to 3.5 ± 0.5 when restricted to a single sector. A similar trend was observed in the 6-sector configuration, where confidence dropped from 5.0 ± 0.0 to 4.1 ± 0.5 under the single-sector condition. These results suggest that users consistently preferred to select multiple sectors, which increased their confidence at identifying the target experimenter. Furthermore, the 6-sector layout led to higher overall confidence, likely because it reduced user concerns about the precision required for accurate selection.
5. LIMITATIONS AND DISCUSSION
Source separation within a sector. Our system currently performs spatial filtering along predefined directional sectors and produces a mixture of all sounds it. This can be undesirable when multiple speakers are close to each other and fall within the same sector. Future extensions of this work could address this limitation in two ways: (1) A low-latency speech separation framework (Gu et al., 2021, 2022; Yuan et al., 2024a) can be applied after directional speech extraction, using permutation-invariant training to separate multiple speakers. (2) More fine-grained separation can be performed in the spatial domain by building on techniques for region-customizable sound extraction (Gu and Luo, 2024) to spatially isolate the target speaker.

### v2

@@ 4.2. Subjective user evaluation of audio quality
While objective metrics like SI-SDR provide valuable quantitative assessment, they cannot fully capture the perceptual quality of speech enhancement systems in real-world scenarios. To evaluate the actual listening experience, we conduct a user study to complement our objective measurements. We randomly select 12 generated audio samples from our test set collected from Room A. The selected audio contains one target speech and 1-2 interfering sources (undesired speech or other sound). Each sample is processed by three systems: the 2-channel microphone array (baseline), 5-channel microphone array, and our proposed SonicSieve. 20 participants (14 males and 6 females) with an average age of 24 years were recruited for the listening test. The evaluation protocol consisted of three steps: (1) listening to the noisy audio mixture, (2) listening to the clean target speaker audio as reference, and (3) listening to the three processed outputs presented in random order to avoid position bias. Participants were asked to rate each processed audio sample on two criteria using 5-point scales. We note that two separate questions (Recommendation, 2003) are used to decouple the assessment of noise removal from the preservation of target speech quality, as aggressive noise suppression can often introduce undesirable artifacts that a single question would fail to capture.
@@
5. LIMITATIONS AND DISCUSSION
Smartphone generalizability. Our system design prioritizes compatibility across different smartphone models through several key considerations:
  • Microphone layout adaptation. Our smartphone application adapts to different device configurations by providing model-specific placement instructions. The UI guides users to position the microstructure optimally based on their device’s microphone locations, ensuring proper alignment between the in-line microphone and the smartphone’s top microphone (Fig. 7).
  • Generalization across smartphone thicknesses. Our evaluation across devices ranging from 10 to 14 mm in thickness (Fig. 13(b)) shows robust performance with only modest degradation for thicker smartphones. These results indicate that the system can generalize across different form factors.
  • Generalization across microphone frequency responses. To enable generalization across different smartphones microphones, device-specific calibration can be performed by measuring the frequency response of the built-in and in-line microphones, and deriving equalization weights to normalize for their effects. This can allow our system to robustly perform directional speech extraction across a range of devices. A complementary approach is the use of established data-driven approaches such as training on diverse device recordings (Li et al., 2017b; Ardila et al., 2019) and targeted data augmentation techniques (Morocutti et al., 2023) which may enable zero-shot generalization without calibration for unseen devices.
@@ 5. LIMITATIONS AND DISCUSSION
Source separation within a sector. Our system currently performs spatial filtering along predefined directional sectors and produces a mixture of all sounds it. This can be undesirable when multiple speakers are close to each other and fall within the same sector. Future extensions of this work could address this limitation in two ways: (1) A low-latency speech separation framework (Gu et al., 2021, 2022; Yuan et al., 2024) can be applied after directional speech extraction, using permutation-invariant training to separate multiple speakers. (2) More fine-grained separation can be performed in the spatial domain by building on techniques for region-customizable sound extraction (Gu and Luo, 2024) to spatially isolate the target speaker.